# Superconducting pairing correlations on a trapped-ion quantum computer

### v2

@@ V Conclusion
Figure S22: Circuit Simulation with Fermioniq’s Ava circuit-DMRG emulator on the light-matter interaction circuits in Fig. 1g. (a) Verification on the exactly solvable $4\times 4$ case, after application of the 2-Trotter step light pulse, before any further relaxation. The exact nearest-neighbour $\Delta_{\eta}$ correlation is obtained by translating the circuit to a 32-qubit Jordan-Wigner circuit and computing the exact statevector result. Compatible with the findings in [haghshenas_digital_2025], (extrapolated) results from circuit-DMRG are reliable when a simulation fidelity $F$ of order 1 can be achieved, and are essentially random for $F\ll 50\%$. (b) Simulation fidelities on the $6\times 6$ circuits on 90 qubits, right after the light pulse (2 steps) and after up to two Hubbard-Trotter relaxation steps (3 and 4 steps), compared with the “intermediate temperature” ($\Delta\theta=2\pi/9$) quench circuit with $s=20$ Ising-Trotter steps in [haghshenas_digital_2025]. (c) Circuit fidelities on 72-qubit circuits that are obtained by deleting all gates on the ancilla qubits, as well as any maximum angle CZZ gates, which are responsible for implementing the fermion-to-qubit encoding. The result is a lower bound on the cost of natively fermionic circuit-DMRG.
@@
We now study the simulability of the circuits with Sparse Pauli Decomposition (SPD). This is a promising simulation technique that has been attracting attention recently [beguvsic2025real, broers2025scalable, lin2025utility, gharibyan2025practical, rudolph2025pauli] and which we have found most practical for the development and testing of quantum circuits in this work. Indeed, it displays a fast and precise convergence for observables with few enough gates in their light-cone for an arbitrary number of qubits (and in general, for circuits with finite magic [dowling2025magic, dowling2025bridging]). This simulation technique works in the Heisenberg picture, and consists in expanding the observable to be measured into strings of Pauli matrices
@@
A generic operator to be decomposed as in (S111) would require $4^{N}$ terms, which becomes intractable for $N\gtrapprox 20$. Starting from a local observable that involves only a few Pauli matrices, we can keep perform the simulation by storing only the non-zero terms in the decomposition (S113). However, the number of gates to keep track of grows exponentially (generically, is multiplied by $2$) after every gate application. The exact bookkeeping of all the non-zero terms ends up thus being unfeasible. As a result, different approximations can be made to get a certain precision on the result. The approximation that we will consider consist in systematically truncating, after every application of gates, all the terms $c_{n}$ that are below some threshold $\epsilon$ in absolute value. Namely, one retains in the sum (S113) only the terms $c_{n}$ such that $|c_{n}|\geq\epsilon$. The complexity of the simulation is related to the non-stabilizerness of the operator, which can be quantified by the operator stabilizer entropy [dowling2025magic]. This simulation technique is particularly appealing for finite-magic circuits and for code testing, as it is not limited by the number of qubits and the circuit geometry. This truncation scheme has been used in the largest-scale SPD numerical simulation to date [broers2025scalable].
@@
In this resource estimation study, we will first focus on the memory requirement, namely on the number of Pauli strings required in the SPD to obtain a certain precision on the observable. We will then discuss an estimation of the actual runtime based on comparison with previous works. We note that Majorana fermion decomposition (see Ref. [miller2025simulation]) instead of Pauli string decompositions are not expected to give an advantage in terms of memory, because within the fermionic encoding there should be a one-to-one correspondence between Pauli strings and Majorana strings.
@@
This is a rather standard metric for estimating the difficulty of the simulation, that has been employed before [gharibyan2025practical]. However, in our case, the circuit has a particular structure. Before the “injection”, the gates at the beginning of the circuit only act on the first $L=36$ qubits, for $90$ qubits in total. This means that in the Heisenberg picture, after the injection is passed, one can keep in the SPD only the strings that have either $I$ or $Z$ Pauli matrices on all the remaining $54$ qubits. This significantly reduces the number of strings to keep track of, but this also modifies the norm, because this projection is non-unitary. Since we would like to know the norm at the end of the circuit, we do not perform this projection. Therefore, for a fairer resource estimate, we will consider the maximal number of strings only before (in the Heisenberg picture) the injection, namely
@@
Figure S25: Relative norm squared as a function of maximal number of strings before injection $\mathcal{S}_{\rm metric}$. Data shown is for the circuits run on hardware on a $6\times 6$ lattice, performing 22 Trotter steps after state preparation (corresponding to the light pulse), 33 Trotter steps (one extra step on top of the light pulse), and 44 Trotter steps (two extra steps). The dashed lines indicate quadratic extrapolations using the last 88 points, and dotted lines linear extrapolations using the last 33 points. The gray shaded area indicates the region that is beyond the largest SPD simulation that has been performed before the present work on a supercomputer [broers2025scalable]. The blue shaded area indicates the region where $90\%$ of the norm is obtained.
@@
How do these numbers translate to actual runtimes? This question is generally difficult to answer due to differences in classical compute hardware and implementation, although in this case we can make some estimates based on the recent parallelised implementation of SPD on the Fugaku supercomputer [broers2025scalable]. In Fig. 3 of that work, wall time per gate is shown for different numbers of Pauli strings $\mathcal{S}$ and parallel processes $M$. Using plot digitisation software [WebPlotDigitizer], we have extracted the scaling of wall time per gate $w$ as a function of $\mathcal{S}$ for a fixed number of parallel processes $M=2^{7}$. While we found the slope $w\sim\mathcal{S}^{1.05}$ to fit the data best, let us make the optimistic assumption that $w\propto\mathcal{S}$ scales linear, that memory is unlimited, and that furthermore the problem can be parallelised perfectly up to $M=2^{22}$ parallel processes. Given that in that work up to 12,288 nodes were used to carry out $2^{18}$ parallel processes, this larger $M$ upper bounds the number of parallel processes that could be run on all 158,976 of Fugaku’s nodes. Under these assumptions, we find a best fit
@@
Figure S26: Estimating the run time for Sparse Pauli Decomposition on Fugaku. (a): Wall time per gate extracted from Fig. 3 in Broers ’25 which refers to Ref. [broers2025scalable] as a function of the number of Pauli strings $\mathcal{S}$. Assuming a perfectly linear scaling of wall time per gate in $\mathcal{S}$ and perfect parallelisation, i.e. inverse linear scaling in the number of parallel processes $M$, the best fit we obtain is $\mathcal{S}/M\times 100$ nanoseconds per gate. (b): Wall time per gate using $M=2^{22}$ parallel processes assuming the scaling extracted in (a) for simulating the 3-step circuit in three different scenarios: A quadratic and linear scenario, as well as the geometric mean, denoted “mid”. We consider the mid scenario to be the most likely of the three and $M^{22}$ to upper bound the maximum number of parallel processes achievable on Fugaku based on Appendix F of [broers2025scalable].
@@
Projected-entangled-pair state (PEPS) methods have become a cornerstone tensor-network framework for simulating quantum many-body systems [cirac2021matrix]. In contrast to MPS, PEPS employs a lattice-based architecture that more naturally and efficiently represents entanglement in higher-dimensional ($D>1$) systems. This enhanced expressiveness, however, comes at the cost of greater computational complexity and the absence of a canonical form for optimal compression. We employ a PEPS specifically adapted to the Octagon lattice geometry with periodic boundary conditions (PBC) and integrate a Belief Propagation (BP) algorithm to efficiently evolve the state [AlAr21].
@@
For comparison, we also report the fidelities [see Eq. (S110)] using MPS simulations, as well as those obtained using matrix-product operator (MPO) Heisenberg evolution. Notably, MPO evolution proceeds backward in time, and its accuracy depends on both the spatial support of the initial (unevolved) operator and the light-cone structure of the circuit. The fidelity of MPO simulations drops rapidly as the light cone of the operator expands, indicating that MPO methods alone are clearly incapable of simulating the full circuit. Similar to both the MPS and SPD approach, linear extrapolation of $\langle O(F)\rangle$ using MPOs has also been considered to extend the limit of finite bond dimension. Empirically, fidelity-based extrapolation of expectation values has been reported to yield results consistent with exact values in the IBM quantum utility experiment [anand2023classical]. However, the expectation value as a function of fidelity is not monotonic. Even as a heuristic method, the fidelity has to be larger than a certain threshold $\mathcal{F}_{\textrm{th}}$ to obtain meaningful signal, either by extrapolation or claiming convergence of the method. Since the MPO and SPD are similar techniques—both track the evolution of observables in the Heisenberg picture and differ only in their truncation schemes—we expect this threshold to be close to $1$, which makes the stand-alone MPO approach practically infeasible and motivates the use of hybrid techniques.
@@
One possible direction to overcome these limitations is a hybrid approach that partially evolves the state in the Schrödinger picture and the operator in the Heisenberg picture, meeting in the middle — a strategy referred to as the mixed representation. For the $4\times 4$ system, one can imagine evolving MPS and MPO and stopping near where their fidelity curves cross to achieve better fidelity. However, given the near-vanishing MPS fidelity for the $6\times 6$ system, one would instead need to evolve PEPS and MPO (or potentially PEPO) and meet at their crossing point to obtain reasonable fidelities. However, measuring observables (in the form of an MPO or PEPO) for PEPS with $\chi>8$ likely requires either a (potentially sign-problematic) sampling technique or a loop-series BP expansion [EvEtAl24]. It is unclear to us whether the observable can be accurately estimated in such a mixed representation with feasible computing resources, but it would be an interesting direction for further study. Finally, we note that all of these methods could also be implemented with an intrinsically fermionic PEPS and/or PEPO representation that avoids additional overhead of simulating the fermion-to-qubit encoding, generally resulting in lower bond dimension for the relevant objects. Reliably establishing the difficulty of each of these potential classical techniques for simulating these circuits requires significant further study and is left for future work.
@@
The second reason for choosing the mixed boundary conditions is that, ultimately, we are interested in the thermodynamic limit $N\to\infty$ where the boundary conditions have no effects anymore. In that respect, periodic, antiperiodic or mixed boundary conditions are only ways of dealing with the finite system, and are physically equally relevant. One can thus choose the boundary condition that will converge the fastest to the thermodynamic limit. It is known that averaging over “twisted” boundary conditions where a phase $e^{i\varphi}$ is given to fermions that wind around the lattice, with $\varphi$ the value to average over, make expectation values of observables converge faster to the thermodynamic limit than periodic boundary conditions [lin2001twist]. In that point of view, the mixed boundary conditions can be seen as a rough approximation of this phase averaging, and is thus expected to give results that are closer to the thermodynamic limit than periodic boundary conditions.
@@
We note that finite-size corrections to the energy density scaling as $L^{-1}$ for open boundaries and $L^{-2}$ for periodic boundaries is a well-known and generic fact in one-dimensional critical systems, see for example [pirvu2012matrix]. In the Octagon encoding, the number of qubits $Q$ is related to the linear system size as $L=\sqrt{2Q/5}$. Therefore, to obtain the same amount of finite-size effects in the two settings, one needs
@@
assuming that the ground state can prepared exactly in each case. We note that, for the free fermion case, the fermionic fast Fourier transform can prepare the ground state reasonably efficiently even without periodic boundary conditions being natively available in the hardware (though reconfigurability does confer a benefit even in this case [maskara2025fastsimulationfermionsreconfigurable]). The quantitative estimate of the advantage of (mixed) periodic qubits shown in Fig. S28c is thus to be seen as a representative example for the simulation of critical states.
@@
For the Hubbard model at half-filling, the quantum Monte Carlo method can provide a numerically exact theory predictions for various physical observables for finite temperatures and the ground state. Such predictions are often used as a metrology tool to determine the temperature of the state prepared in analog quantum simulation [mazurenko_cold-atom_2017, xu_neutral-atom_2025]. Here, we perform the determinant quantum Monte Carlo calculations on a $6\times 6$ lattice with an interaction strength of $U/t=8$ and all four different boundary conditions, i.e., PBC-PBC, PBC-APBC, APBC-PBC, and APBC-APBC, using the QUEST package [varney2009quantum, lee2010quest, chang2015recent]. For the PBC-PBC configuration, we employ the default run script “test”, which exploits the C4 rotational symmetry of the lattice. For all other boundary condition setups, simulations are carried out using the “ggeom” run script, which supports a variety of geometric and boundary configurations. Each simulation consists of 20,000 warmup sweeps followed by 80,000 measurement sweeps with a step size $\tau=0.01$. The mixed boundary condition results reported are the average over the simulation results of the four boundary conditions.
@@
The data supporting this study is available at Zenodo repository with DOI 10.5281/zenodo.17483197 [zenodo].

### v3

@@ V Conclusion
Figure S22: Circuit Simulation with Fermioniq’s Ava circuit-DMRG emulator on the light-matter interaction circuits in Fig. 1g. (a) Verification on the exactly solvable $4\times 4$ case, after application of the 2-Trotter step light pulse, before any further relaxation. The exact nearest-neighbour $\Delta_{\eta}$ correlation is obtained by translating the circuit to a 32-qubit Jordan-Wigner circuit and computing the exact statevector result. Compatible with the findings in [37], (extrapolated) results from circuit-DMRG are reliable when a simulation fidelity $F$ of order 1 can be achieved, and are essentially random for $F\ll 50\%$. (b) Simulation fidelities on the $6\times 6$ circuits on 90 qubits, right after the light pulse (2 steps) and after up to two Hubbard-Trotter relaxation steps (3 and 4 steps), compared with the “intermediate temperature” ($\Delta\theta=2\pi/9$) quench circuit with $s=20$ Ising-Trotter steps in [37]. (c) Circuit fidelities on 72-qubit circuits that are obtained by deleting all gates on the ancilla qubits, as well as any maximum angle CZZ gates, which are responsible for implementing the fermion-to-qubit encoding. The result is a lower bound on the cost of natively fermionic circuit-DMRG.
@@
We now study the simulability of the circuits with Sparse Pauli Decomposition (SPD). This is a promising simulation technique that has been attracting attention recently [11, 16, 50, 33, 68] and which we have found most practical for the development and testing of quantum circuits in this work. Indeed, it displays a fast and precise convergence for observables with few enough gates in their light-cone for an arbitrary number of qubits (and in general, for circuits with finite magic [26, 27]). This simulation technique works in the Heisenberg picture, and consists in expanding the observable to be measured into strings of Pauli matrices
@@
A generic operator to be decomposed as in (S111) would require $4^{N}$ terms, which becomes intractable for $N\gtrapprox 20$. Starting from a local observable that involves only a few Pauli matrices, we can keep perform the simulation by storing only the non-zero terms in the decomposition (S113). However, the number of gates to keep track of grows exponentially (generically, is multiplied by $2$) after every gate application. The exact bookkeeping of all the non-zero terms ends up thus being unfeasible. As a result, different approximations can be made to get a certain precision on the result. The approximation that we will consider consist in systematically truncating, after every application of gates, all the terms $c_{n}$ that are below some threshold $\epsilon$ in absolute value. Namely, one retains in the sum (S113) only the terms $c_{n}$ such that $|c_{n}|\geq\epsilon$. The complexity of the simulation is related to the non-stabilizerness of the operator, which can be quantified by the operator stabilizer entropy [26]. This simulation technique is particularly appealing for finite-magic circuits and for code testing, as it is not limited by the number of qubits and the circuit geometry. This truncation scheme has been used in the largest-scale SPD numerical simulation to date [16].
@@
In this resource estimation study, we will first focus on the memory requirement, namely on the number of Pauli strings required in the SPD to obtain a certain precision on the observable. We will then discuss an estimation of the actual runtime based on comparison with previous works. We note that Majorana fermion decomposition (see Ref. [57]) instead of Pauli string decompositions are not expected to give an advantage in terms of memory, because within the fermionic encoding there should be a one-to-one correspondence between Pauli strings and Majorana strings.
@@
This is a rather standard metric for estimating the difficulty of the simulation, that has been employed before [33]. However, in our case, the circuit has a particular structure. Before the “injection”, the gates at the beginning of the circuit only act on the first $L=36$ qubits, for $90$ qubits in total. This means that in the Heisenberg picture, after the injection is passed, one can keep in the SPD only the strings that have either $I$ or $Z$ Pauli matrices on all the remaining $54$ qubits. This significantly reduces the number of strings to keep track of, but this also modifies the norm, because this projection is non-unitary. Since we would like to know the norm at the end of the circuit, we do not perform this projection. Therefore, for a fairer resource estimate, we will consider the maximal number of strings only before (in the Heisenberg picture) the injection, namely
@@
Figure S25: Relative norm squared as a function of maximal number of strings before injection $\mathcal{S}_{\rm metric}$. Data shown is for the circuits run on hardware on a $6\times 6$ lattice, performing 22 Trotter steps after state preparation (corresponding to the light pulse), 33 Trotter steps (one extra step on top of the light pulse), and 44 Trotter steps (two extra steps). The dashed lines indicate quadratic extrapolations using the last 88 points, and dotted lines linear extrapolations using the last 33 points. The gray shaded area indicates the region that is beyond the largest SPD simulation that has been performed before the present work on a supercomputer [16]. The blue shaded area indicates the region where $90\%$ of the norm is obtained.
@@
How do these numbers translate to actual runtimes? This question is generally difficult to answer due to differences in classical compute hardware and implementation, although in this case we can make some estimates based on the recent parallelised implementation of SPD on the Fugaku supercomputer [16]. In Fig. 3 of that work, wall time per gate is shown for different numbers of Pauli strings $\mathcal{S}$ and parallel processes $M$. Using plot digitisation software [67], we have extracted the scaling of wall time per gate $w$ as a function of $\mathcal{S}$ for a fixed number of parallel processes $M=2^{7}$. While we found the slope $w\sim\mathcal{S}^{1.05}$ to fit the data best, let us make the optimistic assumption that $w\propto\mathcal{S}$ scales linear, that memory is unlimited, and that furthermore the problem can be parallelised perfectly up to $M=2^{22}$ parallel processes. Given that in that work up to 12,288 nodes were used to carry out $2^{18}$ parallel processes, this larger $M$ upper bounds the number of parallel processes that could be run on all 158,976 of Fugaku’s nodes. Under these assumptions, we find a best fit
@@
Figure S26: Estimating the run time for Sparse Pauli Decomposition on Fugaku. (a): Wall time per gate extracted from Fig. 3 in Broers ’25 which refers to Ref. [16] as a function of the number of Pauli strings $\mathcal{S}$. Assuming a perfectly linear scaling of wall time per gate in $\mathcal{S}$ and perfect parallelisation, i.e. inverse linear scaling in the number of parallel processes $M$, the best fit we obtain is $\mathcal{S}/M\times 100$ nanoseconds per gate. (b): Wall time per gate using $M=2^{22}$ parallel processes assuming the scaling extracted in (a) for simulating the 3-step circuit in three different scenarios: A quadratic and linear scenario, as well as the geometric mean, denoted “mid”. We consider the mid scenario to be the most likely of the three and $M^{22}$ to upper bound the maximum number of parallel processes achievable on Fugaku based on Appendix F of [16].
@@
Projected-entangled-pair state (PEPS) methods have become a cornerstone tensor-network framework for simulating quantum many-body systems [21]. In contrast to MPS, PEPS employs a lattice-based architecture that more naturally and efficiently represents entanglement in higher-dimensional ($D>1$) systems. This enhanced expressiveness, however, comes at the cost of greater computational complexity and the absence of a canonical form for optimal compression. We employ a PEPS specifically adapted to the Octagon lattice geometry with periodic boundary conditions (PBC) and integrate a Belief Propagation (BP) algorithm to efficiently evolve the state [2].
@@
For comparison, we also report the fidelities [see Eq. (S110)] using MPS simulations, as well as those obtained using matrix-product operator (MPO) Heisenberg evolution. Notably, MPO evolution proceeds backward in time, and its accuracy depends on both the spatial support of the initial (unevolved) operator and the light-cone structure of the circuit. The fidelity of MPO simulations drops rapidly as the light cone of the operator expands, indicating that MPO methods alone are clearly incapable of simulating the full circuit. Similar to both the MPS and SPD approach, linear extrapolation of $\langle O(F)\rangle$ using MPOs has also been considered to extend the limit of finite bond dimension. Empirically, fidelity-based extrapolation of expectation values has been reported to yield results consistent with exact values in the IBM quantum utility experiment [3]. However, the expectation value as a function of fidelity is not monotonic. Even as a heuristic method, the fidelity has to be larger than a certain threshold $\mathcal{F}_{\textrm{th}}$ to obtain meaningful signal, either by extrapolation or claiming convergence of the method. Since the MPO and SPD are similar techniques—both track the evolution of observables in the Heisenberg picture and differ only in their truncation schemes—we expect this threshold to be close to $1$, which makes the stand-alone MPO approach practically infeasible and motivates the use of hybrid techniques.
@@
One possible direction to overcome these limitations is a hybrid approach that partially evolves the state in the Schrödinger picture and the operator in the Heisenberg picture, meeting in the middle — a strategy referred to as the mixed representation. For the $4\times 4$ system, one can imagine evolving MPS and MPO and stopping near where their fidelity curves cross to achieve better fidelity. However, given the near-vanishing MPS fidelity for the $6\times 6$ system, one would instead need to evolve PEPS and MPO (or potentially PEPO) and meet at their crossing point to obtain reasonable fidelities. However, measuring observables (in the form of an MPO or PEPO) for PEPS with $\chi>8$ likely requires either a (potentially sign-problematic) sampling technique or a loop-series BP expansion [28]. It is unclear to us whether the observable can be accurately estimated in such a mixed representation with feasible computing resources, but it would be an interesting direction for further study. Finally, we note that all of these methods could also be implemented with an intrinsically fermionic PEPS and/or PEPO representation that avoids additional overhead of simulating the fermion-to-qubit encoding, generally resulting in lower bond dimension for the relevant objects. Reliably establishing the difficulty of each of these potential classical techniques for simulating these circuits requires significant further study and is left for future work.
@@
The second reason for choosing the mixed boundary conditions is that, ultimately, we are interested in the thermodynamic limit $N\to\infty$ where the boundary conditions have no effects anymore. In that respect, periodic, antiperiodic or mixed boundary conditions are only ways of dealing with the finite system, and are physically equally relevant. One can thus choose the boundary condition that will converge the fastest to the thermodynamic limit. It is known that averaging over “twisted” boundary conditions where a phase $e^{i\varphi}$ is given to fermions that wind around the lattice, with $\varphi$ the value to average over, make expectation values of observables converge faster to the thermodynamic limit than periodic boundary conditions [48]. In that point of view, the mixed boundary conditions can be seen as a rough approximation of this phase averaging, and is thus expected to give results that are closer to the thermodynamic limit than periodic boundary conditions.
@@
We note that finite-size corrections to the energy density scaling as $L^{-1}$ for open boundaries and $L^{-2}$ for periodic boundaries is a well-known and generic fact in one-dimensional critical systems, see for example [61]. In the Octagon encoding, the number of qubits $Q$ is related to the linear system size as $L=\sqrt{2Q/5}$. Therefore, to obtain the same amount of finite-size effects in the two settings, one needs
@@
assuming that the ground state can prepared exactly in each case. We note that, for the free fermion case, the fermionic fast Fourier transform can prepare the ground state reasonably efficiently even without periodic boundary conditions being natively available in the hardware (though reconfigurability does confer a benefit even in this case [54]). The quantitative estimate of the advantage of (mixed) periodic qubits shown in Fig. S28c is thus to be seen as a representative example for the simulation of critical states.
@@
For the Hubbard model at half-filling, the quantum Monte Carlo method can provide a numerically exact theory predictions for various physical observables for finite temperatures and the ground state. Such predictions are often used as a metrology tool to determine the temperature of the state prepared in analog quantum simulation [56, 84]. Here, we perform the determinant quantum Monte Carlo calculations on a $6\times 6$ lattice with an interaction strength of $U/t=8$ and all four different boundary conditions, i.e., PBC-PBC, PBC-APBC, APBC-PBC, and APBC-APBC, using the QUEST package [79, 47, 19]. For the PBC-PBC configuration, we employ the default run script “test”, which exploits the C4 rotational symmetry of the lattice. For all other boundary condition setups, simulations are carried out using the “ggeom” run script, which supports a variety of geometric and boundary configurations. Each simulation consists of 20,000 warmup sweeps followed by 80,000 measurement sweeps with a step size $\tau=0.01$. The mixed boundary condition results reported are the average over the simulation results of the four boundary conditions.
@@
The data supporting this study is available at Zenodo repository with DOI 10.5281/zenodo.17483197 [36].